# Can LLM Generate Culturally Relevant Commonsense QA Data? Case Study in Indonesian and Sundanese

### v2

@@ 1 Introduction
Another common method for constructing non-English datasets is by utilizing a machine translation model. Although more scalable, this method cannot be straightforwardly applied due to the contextual irrelevancy of the data, primarily due to geographical differences (which could also influence cultural differences). For example, many English CommonsenseQA questions Talmor et al. (2019) include concepts such as snow or any winter sports, which are irrelevant in Indonesia due to its tropical climate with only two seasons. This dataset also often includes English-centric names and locations, primarily limited to the US, which are not considered “commonsense” for Indonesians.
@@ 3.1 Commonsense QA: Definition and Scope
In our dataset, we focus on building questions that probe common or cultural knowledge in daily life within Indonesian and Sundanese contexts. Each data point is a triple of concept, question, and options, with one correct answer. To enhance the relevance of the dataset to local contexts and reduce Western cultural bias, we not only utilized existing question concepts from the English CommonsenseQA Talmor et al. (2019) but also manually created novel question concepts. We cover five categories: (1) culinary, (2) place, (3) culture, (4) history, and (5) activity, with 150 unique question concepts for each category, aiming for broad domain and knowledge coverage.
@@ 4 Data Generation Methods
To investigate whether LLMs can generate culturally relevant commonsense QA data in Indonesian and Sundanese, we construct a dataset using various methods with LLMs as data generators. In addition, we employ humans to generate data for comparison. As illustrated in Figure 1, we apply three dataset generation methods: (1) Automatic Data Adaptation (LLM_Adapt), where we leverage LLMs to automatically adapt English CommonsenseQA data to our target languages; (2) Manual Data Generation (Human_Gen), where we ask native-speaker human annotators to manually construct the dataset; and (3) Automatic Data Generation (LLM_Gen), where we use LLMs to generate data based on the list of categories and concepts used in Human_Gen. As shown in Table 1, our final dataset contains 8,953 QA pairs in total (4,416 for Indonesian and 4,537 for Sundanese).
@@
Table 1: Statistics of our generated Indonesian and Sundanese CommonsenseQA dataset.
@@ 4.1 Automatic Data Adaptation
We build the first LLM-generated data by adapting the CommonsenseQA dataset Talmor et al. (2019) to make it culturally relevant to Indonesian and Sundanese. First, we remove data that contains ambiguous options and offensive keywords. Next step involves assessing three elements: concept, name, and location, in order to sample the data that will be adapted. We take the question concepts from the existing CommonsenseQA data. As for the name and location extraction, we use Stanford’s Stanza NLP library Qi et al. (2020) and ConceptNet Speer et al. (2017) API. We use GPT-3.5 Turbo results ensembled from five different prompts to determine the relevance of the data based on the three elements. Data that are considered irrelevant in at least one of the three elements are selected to be adapted. The examples are shown in Table 11.
The next step is to transform the selected irrelevant data. First, we prompt GPT-4 Turbo [^2] Our initial experiment (Appendix A.2) reveals that GPT-4 Turbo significantly outperforms Merak-v4, an open Indonesian LLM, leading us to select GPT-4 Turbo for our work. to rephrase the sampled questions and options to align with Indonesian cultures. Subsequently, for data flagged with Western-centric names, a second rephrasing is conducted using the GPT-3.5 Turbo to replace all person names. The fully rephrased and adapted data are then translated from Indonesian to Sundanese using Google Translation API. Finally, to ensure the dataset quality, we filter the data by examining the appearance of concepts in the questions and checking Sundanese-Indonesian back-translation with a similarity threshold of 90%.
@@ 4.2 Manual Data Generation
To build the human-generated dataset, we recruited 12 experienced annotators from diverse regions across Java and Bali, as detailed in Table 10. All are native speakers of Indonesian (and Sundanese), with a minimum of 15 years residing in their particular target language regions. Among them, 5 annotators have a degree in linguistics. These criteria ensure they possess both language proficiency and cultural familiarity. As for the data collection process, we have two main phases: (1) creating commonsense question-answer pairs and (2) answering commonsense questions. We also perform quality control to ensure the data quality. [^3] More details on annotators’ demographic information are presented in Appendix B. The complete annotation guideline is outlined in Appendix C.
Creating Question-Answer Pairs
We first instruct the annotators to create commonsense question-answer pairs based on the given category and question concept that we had newly created (§3.1). We also ask them to rely on their existing knowledge when making questions and avoid using internet search or LLM as much as possible. Two annotators use the same 50 question concepts from each category to generate the questions, options, and answers. This process produces 1,500 QA pairs for each language in total.
@@ 4.3 Automatic Data Generation
For generating the second type of LLM-generated data, rather than adapting questions from the English dataset (§4.1), we use the same set of categories and question concepts as the human-generated data (§4.2). We also utilize GPT-4 Turbo and instruct it to generate questions, options, and answers. To ensure the generated dataset aligns closely with the intended cultural context, we explicitly incorporated the categories and question concepts in the prompt. Additionally, we instruct the model to strictly include the question concepts in the generated questions. We adopt a batching approach to streamline the data generation process, providing the model with a maximum of 5 distinct question concepts from the same categories in one API call. To maintain dataset quality and uniqueness, we filter out duplicate entries and questions that do not explicitly contain the question concepts.
@@ 5.1 LLM-Generated Data
To evaluate the quality of our LLM-generated data, we manually reviewed all 158 samples from the LLM_Adapt test set. Based on the evaluation, we then calculate the accuracy of the generated concepts, questions, and options. Furthermore, we evaluate 300 randomly selected samples from the LLM_Gen using the same procedure.
5.1.1 Concept Analysis
Concept Quality
In the English to Indonesian concept adaptation, 95.57% are correct, including 16 out of 19 concepts needing major adaptation for cultural relevance, such as ‘snow’ to ‘hujan abu vulkanik’ (volcanic ashfall), showing that the LLM has some basic knowledge to adapt English concepts, although most adaptations primarily involving exact translations without major alteration of the concept (Table 2). For Indonesian to Sundanese, the correct adaptation drops to 77.22%, reflecting weak machine translation performance in Sundanese (Table 8).
@@ 5.1.1 Concept Analysis
Despite the high accuracy of concept adaptation, as shown in Figure 2, the adapted concepts are skewed towards ‘komodo,’ indicating a bias towards a specific entity within a category. This could be due to the model’s limited knowledge or the absence of a direct equivalent for the original English concept in Indonesian, leading it to default to one standard concept. This finding highlights the limitations of adaptation and translation methods, and to improve dataset diversity and coverage, we need to manually develop concepts across various categories.
@@ 5.1.2 Question Analysis
To evaluate the quality of the generated questions, we apply a strict criterion: any errors, even minor, are marked as incorrect. As shown in Table 3, the Indonesian datasets show a high percentage of error-free questions, ranging between 68–75%. However, for Sundanese, the accuracy decreases significantly. In particular, the weak performance of the MT system in Sundanese is evident from its very low accuracy in LLM_Adapt, suggesting synthetic data adaptation/translation from English is not an optimal method for low-resource languages. In contrast, the LLM_Gen dataset, which involves generating synthetic data directly in the target language, shows a significantly higher number of correct questions, particularly for Sundanese. This indicates that direct generation in the target language is a more promising approach than adaptation or translation.
@@
We manually reviewed the questions to identify common errors, detailed in Table 5.1.3. We observe that most of the errors in LLM_Adapt come from translation errors. For Indonesian LLM_Gen, despite lower question generation accuracy, most of the errors are actually minor typos. In Sundanese LLM_Gen, however, the predominant issues relate to sentence fluency. This indicates that the model is still struggling to generate fluent sentences in Sundanese, where it occasionally uses phrases or words in Indonesian instead of Sundanese. While Indonesian and Sundanese share many linguistic features, they differ significantly in morphological features (§3.2). This characteristic, along with the smaller amount of Sundanese data used for training LLMs, results in the models being less exposed to Sundanese words, suffixes, or prefixes compared to Indonesian. This leads to the occasional appearance of Indonesian words in Sundanese data, even though the sentences may be grammatically correct.
@@ 5.1.3 Options and Answer Analysis
Table 3: Question generation accuracy of LLM-generated datasets. The accuracy reflects the percentage of a ‘clean’ question without any major or minor errors.
@@
Table 5: Options generation accuracy of LLM-generated datasets, measured by the % of clean/error-free options.
@@
We also evaluate the options in a similar method to the evaluation of the question. From Table 5, we observe that LLM_Gen generates higher quality options compared to LLM_Adapt. However, there is still a significant gap in performance between Indonesian and Sundanese, once again highlighting the performance discrepancy between medium- and lower-resource languages.
@@
As detailed in Table 5.1.3, for LLM_Adapt, the most common issues are minor typos and mechanical errors, particularly with capitalization (e.g., "indonesia" instead of "Indonesia"). For Sundanese, the errors are mainly due to major translation errors. As for LLM_Gen, the most common errors involve the presence of invalid options, particularly when no correct answers are among the options. For Sundanese, alongside translation errors, the model also produces some invalid options. However, it is important to note that the proportion of invalid options is very low for both languages, indicating that the model is still capable of generating a large number of valid options.
@@
Upon some sample-level analysis, we find that given the same set of categories and question concepts, humans generate more token variations that are not produced by LLMs, such as some unique terms like kalis or cimol. [^4] In the context of culinary, ‘kalis’ means a state of dough that is well-kneaded and ready to be processed further. ‘Cimol’ is an Indonesian street food made from tapioca flour. Although LLM can still generate some valid concepts, it tends to ask more general questions or popular concepts. For instance, given ‘kerupuk’ (crackers) concept, human annotators can formulate questions tied to their cultural background, such as asking about ‘kerupuk rambak’ (rambak crackers). In contrast, LLM tends to ask more general (but still relevant) questions, like "What is the common primary ingredient of crackers?" More examples are shown in Table 12.
@@ 6.1 Experiment Setup
We use three prompt variations for all models (see Appendix E). For a fair comparison between open and proprietary LLMs, we extract the answer key from the text generation result instead of the next token probability, using a rule-based and regex. Evaluations are performed on RTX A6000 48GB. We use accuracy as the evaluation metric.
Figure 3: Overall performance of LLMs on our combined test set.
Figure 4: LLMs’ performance on LLM_Gen vs. Human_Gen dataset in Indonesian and Sundanese.
@@ 6.2 Overall Performance
We also observe a substantial gap between Indonesian and Sundanese (10–20% accuracy drop), suggesting that current LLMs struggle to answer our Sundanese commonsense questions, even in multiple-choice settings. The huge gap, particularly in Merak-v4, highlights the limitations of training LLMs solely on Indonesian texts, as it does not ensure transferable performance across other local languages, especially due to the morphological differences between the languages. Including Sundanese texts in the training data, as demonstrated by improved performance in the Aya model, has proven effective in narrowing the performance gap between Indonesian and Sundanese.
@@
To assess LLMs’ ability to answer LLM and human-generated data, we compare their performances on our LLM_Gen and Human_Gen datasets. Figure 4 shows that LLMs generally perform higher on LLM_Gen, especially on MistalOrca, which is trained on English-centric data. This indicates that the model may possess some knowledge to answer LLM_Gen questions, even though it was not specifically trained on Indonesian or Sundanese data. Still, it is important to note that even in LLM_Gen, MistalOrca’s average score is still lower compared to LLMs that include Indonesian texts in their training.
Interestingly, as the model gets stronger, we observe that the gap tends to narrow, especially in GPT models. For Aya, despite its relatively stronger performance compared to other open-source LLMs, its performance is significantly dropped in Human_Gen data. This decline is likely due to Aya being a ‘translation-heavy’ model, with a large proportion of its training data derived from translations rather than human annotations.

### v3

@@ 1 Introduction
Another common method for constructing non-English datasets is by using machine translation. Although more scalable, this method cannot be straightforwardly applied due to the contextual irrelevancy of the data, primarily due to geographical differences (which could also influence cultural differences). For example, many English CommonsenseQA questions Talmor et al. (2019) include concepts such as snow or any winter sports, which are irrelevant in Indonesia due to its tropical climate with only two seasons. This dataset also often includes English-centric names and locations, primarily limited to the US, which are not considered “commonsense” for Indonesians.
@@ 3.1 Commonsense QA: Definition and Scope
In our dataset, we focus on building questions that probe common or cultural knowledge in daily life within Indonesian and Sundanese contexts. Each data point is a triple of concept, question, and options, with one correct answer. To enhance the relevance of the dataset to local contexts and reduce Western cultural bias, we incorporate both existing question concepts from the English CommonsenseQA Talmor et al. (2019) and novel question concepts that we manually created.
Unlike English CommonsenseQA, our dataset includes question category metadata, covering five categories: culinary, place, culture, history, and activity. Each category has 150 unique question concepts, ensuring broad domain and knowledge coverage. [^2] More details on how we curate our categories and concepts can be seen in Appendix B. Another key difference is that English CommonsenseQA concepts are sourced from a knowledge base, and they divide the concept sets into questions and options. In contrast, our dataset provides only a question concept and its category as input. The options, including the correct answer, are entirely generated by the data creator.
@@ 4 Data Generation Methods
To investigate whether LLMs can generate culturally relevant commonsense QA data in Indonesian and Sundanese, we build a dataset using various methods with LLMs as data generators. We also employ humans to generate data for comparison.
As illustrated in Figure 1, we apply three dataset generation methods: (1) Automatic Data Adaptation (LLM_Adapt), where we leverage LLMs to automatically adapt English CommonsenseQA data to our target languages; (2) Manual Data Generation (Human_Gen), where we ask native-speaker human annotators to manually construct the dataset; and (3) Automatic Data Generation (LLM_Gen), where we use LLMs to generate data based on the list of categories and concepts used in Human_Gen. Table 1 shows the statistics of our final dataset, containing a total of 8,953 QA pairs (4,416 Indonesian and 4,537 Sundanese).
@@
Table 1: Statistics of our generated Indonesian and Sundanese CommonsenseQA dataset. We retained the original English CommonsenseQA splits in LLM_Adapt to avoid data contamination.
@@ 4.1 Automatic Data Adaptation
We build the first LLM-generated data by adapting English CommonsenseQA dataset Talmor et al. (2019) to make it culturally relevant to Indonesian and Sundanese.
Data Selection
To select the data to be adapted, we first remove data containing offensive keywords and those with duplicate or similar options in different tenses. Unlike English, Indonesian and Sundanese do not change form to indicate tenses (e.g., ‘passed’ and ‘passing’ both translate to ‘lulus’).
Then, we sample the data by assessing three elements: concept, name, and location. Data that are considered irrelevant in at least one of the three elements are selected to be adapted. We take question concepts from the existing CommonsenseQA data and use Stanza Qi et al. (2020) and ConceptNet API [^3] https://github.com/commonsense/conceptnet5/wiki/API for name and location extraction.
[…]
QA Pairs Adaptation
The next step is to transform the selected irrelevant data. First, we prompt GPT-4 Turbo [^5] Our initial experiment (Appendix A.2) reveals that GPT-4 Turbo significantly outperforms Merak-v4, an open Indonesian LLM, leading us to select GPT-4 Turbo for our work. to rephrase the sampled questions and options to align with Indonesian cultures. Subsequently, for data flagged with Western-centric names, we use GPT-3.5 Turbo to replace all person names. The fully adapted data are then translated from Indonesian to Sundanese using Google Translation API. We choose machine translation due to the unreliability of direct adaptation, as the concepts are sourced from English data. Our preliminary run on eng to sun adaptation shows that 90% of the samples contained errors, with half due to hallucinations. This occurred less frequently in eng to ind adaptation. For example, for the ‘bald eagle’ concept, GPT-4 Turbo generated ‘elang Jawa’ (Javan hawk-eagle) in Indonesian, but it generated the non-existent concept ‘garuda puspa’ (literally translates to eagle flower) in Sundanese. Although not ideal, translation from the adapted Indonesian QA pairs generally provides better data quality.
Data Filtering
Finally, we eliminate low-quality data by removing instances where the concepts do not appear in the questions. We also utilize a back-translation method to filter out poor translations of Indonesian (ind) to Sundanese (sun) data. We discard QA pairs if the similarity between the original Indonesian (ind) and the back-translated version (ind’) is less than 0.9. We measure similarity using an embedding-based metric with multilingual MiniLM Reimers and Gurevych (2019) and LaBSE Feng et al. (2022) models, and keep the data if either similarity score is above 0.9.
@@ 4.2 Manual Data Generation
To build the human-generated dataset, we recruited 12 experienced annotators from diverse regions across Java and Bali. All are native speakers of Indonesian (and Sundanese), with a minimum of 15 years residing in their particular target language regions. Among them, 5 annotators have a degree in linguistics. These criteria ensure they possess both language proficiency and cultural familiarity. As for the data collection process, we have two main phases: (1) creating commonsense question-answer pairs and (2) answering commonsense questions. We also perform quality control to ensure the data quality. [^6] More details on annotators’ demographic information and annotation guidelines are presented in Appendix C and D.
QA Pairs Creation
We first instruct the annotators to create commonsense question-answer pairs based on the given category and question concept that we had newly created (§3.1). We also ask them to rely on their existing knowledge when making questions and avoid using internet search or LLMs as much as possible. Each concept is annotated by 2 annotators, with each annotator covering all 5 categories to ensure consistency. Since we have 150 unique concepts per category and 6 annotators per language, each annotator is assigned to create QA pairs for 50 concepts per category. This results in 300 QA pairs per category (150 concepts $\times$× 2 annotators) and a total of 1,500 QA pairs per language (300 QA pairs $\times$× 5 categories).
@@ 4.3 Automatic Data Generation
For generating the second type of LLM-generated data, rather than adapting questions from the English dataset (§4.1), we use the same set of categories and question concepts as the human-generated data (§4.2). We also utilize GPT-4 Turbo and instruct it to generate questions, options, and answers. To ensure the generated dataset aligns closely with the intended cultural context, we explicitly incorporated the categories and question concepts in the prompt. Additionally, we instruct the model to strictly include the question concepts in the generated questions.
We adopt a batching approach to streamline the data generation process, providing the model with a maximum of 5 distinct question concepts from the same categories in one API call. We chose 5 concepts because this number is optimal; larger batches would exceed the API’s maximum length, while smaller batches would result in longer processing times.
To maintain dataset quality and uniqueness, we filter out duplicate entries and questions that do not explicitly contain the question concepts.
@@ 5.1 LLM-Generated Data
To evaluate the quality of our LLM-generated data, we manually reviewed all 158 samples from the LLM_Adapt test set (see Table 1). Based on the evaluation, we then calculate the accuracy of the generated concepts, questions, and options. We also evaluate 300 randomly selected samples from the LLM_Gen using the same procedure.
@@
Status Num (%) Concept Example
Orig (ind) Modified (sun)
Correct 122 (77.22%)
hujan abu vulkanik
(volcanic ashfall)
hujan lebu vulkanik
(volcanic ashfall)
menara (tower) munara (tower)
Wrong 36 (22.78%)
cicak
(house gecko)
kadal imah
(house lizard)
klinik gigi
(dental clinic)
klinik dental
(dental clinic)
Table 3: ind to sun concept adaptation result.
5.1.1 Concept Analysis
Concept Quality
To evaluate the quality of the adapted concepts, we regard the concept as correct if it is a real, existing concept (not a hallucination) and relevant to the Indonesian or Sundanese context. As shown in Table 2, 95.57% of English to Indonesian adaptations are correct, including 16 out of 19 concepts needing major adaptation, such as ‘snow’ to ‘hujan abu vulkanik’ (volcanic ashfall). This result indicates that the LLM can adequately adapt some English concepts to Indonesian, although most involve direct translations without major alterations. However, for Indonesian to Sundanese (Table 3), the correct adaptation drops to 77.22%, reflecting weak machine translation (MT) performance in Sundanese.
@@ 5.1.1 Concept Analysis
Despite the high accuracy of concept adaptation, as shown in Figure 2, the adapted concepts are skewed towards ‘komodo,’ indicating a bias towards a specific entity within a category. This could be due to the model being trained on data with insufficient knowledge or the absence of a direct equivalent of the English concept in Indonesian, leading it to default to one standard concept. This finding highlights the limitations of concept adaptation from existing English data. To improve dataset diversity and coverage, manual concept creation in the target language is needed.
@@ 5.1.2 Question Analysis
To evaluate the quality of the generated questions, we apply a strict criterion: any errors, even minor, are marked as incorrect. As shown in Table 4, the Indonesian datasets have a high percentage of error-free questions, ranging between 68–75%. However, for Sundanese, the accuracy decreases significantly. In particular, the weak performance of the MT system in Sundanese is evident from its very low accuracy in LLM_Adapt, suggesting synthetic data adaptation/translation from English is not an optimal method for low-resource languages. In contrast, the LLM_Gen dataset, which involves generating synthetic data directly in the target language, shows a significantly higher number of correct questions, particularly for Sundanese. This indicates that direct generation in the target language is a more promising approach than adaptation or translation.
@@
We manually reviewed the questions to identify common errors, detailed in Table 5.1.3. We observe that most of the errors in LLM_Adapt come from translation errors. For Indonesian LLM_Gen, despite lower question generation accuracy, most of the errors are actually minor typos. In Sundanese LLM_Gen, however, the predominant issues relate to sentence fluency, indicating ongoing challenges in automatically generating smooth and fluent sentences in Sundanese. We also found that the model occasionally outputs wrong languages, particularly using Indonesian phrases or words instead of Sundanese. While Indonesian and Sundanese share many linguistic features, they differ significantly in morphological features (§3.2). This difference, along with the smaller amount of Sundanese data used for training LLMs, results in the models being less exposed to Sundanese words, suffixes, or prefixes compared to Indonesian. This leads to the occasional appearance of Indonesian words in generated Sundanese data, even though the sentences may be grammatically correct.
@@ 5.1.3 Options and Answer Analysis
Table 4: Question generation acc. of LLM-generated datasets, measured by the % of error-free questions.
@@
Table 6: Options generation acc. of LLM-generated datasets, measured by the % of error-free options.
@@
We also assess the options quality using a similar method to the question evaluation. From Table 6, we find LLM_Gen generates higher quality options compared to LLM_Adapt. However, there is still a significant gap in performance between Indonesian and Sundanese, once again highlighting the performance discrepancy between medium- and lower-resource languages.
@@
As detailed in Table 5.1.3, for LLM_Adapt, the most common issues are minor typos and mechanical errors, particularly with capitalization (e.g., "indonesia" instead of "Indonesia"). For Sundanese, the errors are mainly due to major translation errors. As for LLM_Gen, the most common errors involve the presence of invalid options, particularly when no correct answers are among the options. For Sundanese, alongside invalid options, the model also produces some options in the wrong language. Still, it is important to note that the proportion of invalid options is very low for both languages, indicating that the model generates a substantial number of valid options.
@@
Upon sample-level analysis, we find that given the same set of categories and question concepts, humans generate more token variations that are not produced by LLMs, such as some unique terms like kalis or cimol. [^7] In the context of culinary, ‘kalis’ means a state of dough that is well-kneaded and ready to be processed further. ‘Cimol’ is an Indonesian street food made from tapioca flour. Although LLM can still generate some valid data, it tends to return more popular concepts or general questions. For instance, given ‘kerupuk’ (crackers) concept, human annotators can create questions tied to their cultural background, such as asking about ‘kerupuk rambak’ (rambak crackers). In contrast, LLM tends to give more general (but still relevant) questions, like "What is the common primary ingredient of crackers?" More examples are shown in Table 12 in the Appendix.
@@ 6.1 Experiment Setup
We use three prompt variations for all models. [^8] Please refer to Appendix F for the prompts details. For a fair comparison between open and proprietary LLMs, we extract the answer key from the text generation result instead of the next token probability, using a rule-based and regex. Evaluations are performed on RTX A6000 48GB. We use accuracy as the evaluation metric.
@@ 6.2 Overall Performance
We also observe a substantial gap between Indonesian and Sundanese (10–20% accuracy drop), suggesting that current LLMs struggle with Sundanese questions, even in multiple-choice settings. This gap, particularly in Merak-v4, highlights the limitations of training LLMs solely on Indonesian texts, which does not ensure transferable performance across other local languages due to morphological differences. Including Sundanese texts in the training data, as shown by the improved performance in Aya, effectively narrows the performance gap between Indonesian and Sundanese.
@@
Figure 4: LLMs’ performance on LLM_Gen vs. Human_Gen in Indonesian and Sundanese. We combined data points from both languages for visualization, with lower quartiles typically representing Sundanese data.
To assess LLMs’ ability to answer LLM and human-generated data, we compare their performances on our LLM_Gen and Human_Gen datasets. Figure 4 shows that LLMs generally perform higher on LLM_Gen, especially on MistalOrca, which is trained on English-centric data. This suggests the model can handle LLM_Gen questions despite not being specifically trained on Indonesian or Sundanese data. Still, it is important to note that MistalOrca’s average score on LLM_Gen is still lower than other LLMs that include Indonesian texts in their training.
Interestingly, the performance gap tends to narrow as models improve, particularly for GPT models. For Aya, despite relatively strong performance compared to other open-source LLMs, shows significantly dropped performance in Human_Gen data. This decline is likely due to Aya being a ‘translation-heavy’ model Üstün et al. (2024), with a large proportion of its training data derived from translations rather than human annotations.